xfcghvb:

One week, Isabelle bought 2 bags of apples and 3 bags of oranges. The next week, she bought 1 bag of apples and 2 bags of oranges. If a represents the cost of each bag of apples and r represents the cost of each bag of oranges, which expression could Isabelle use to calculate the difference between the amounts she spent each week?

Shadow:

\[2a + 3r = w_{1}\] \[1a + 2r = w_{2}\]

Shadow:

1a is actually simplified to just a

Shadow:

Written out, that's exactly what the paragraph says. Week one is equal to two apples plus three oranges. Week two is equal to one apple plus two oranges.

xfcghvb:

how do you get the answer tho i dont get it

xfcghvb:

Which expression is equivalent to (–3y – x) – (5y – 8x)? –8y – 8x –8y + 7x 2y – 7x 8y + 8x

Shadow:

The difference is week one minus week two Or: \[w_{1} - w_{2}\] \[(2a + 3r) - (a + 2r) = difference\]

xfcghvb:

thank you